OpenStudy (xmoses1):

If x varies inversely with y and x = 4 when y = 8, find x when y = 16. 32 8 2 4

OpenStudy (anonymous):

So when y =8 and x =4, We get, 8=k/4 32=k Since k is a constant it will remain the same for the statement 'y varies inversely as x'. Now when x = 16, y= 32/16 y=2

OpenStudy (xmoses1):

Thank you! :)